# AIM 2024 Challenge on Compressed Video Quality Assessment: Methods and Results

### v1

@@ 1 Introduction
This challenge is one of the AIM 2024 Workshop challenges: Mobile Real-Time Video Super-Resolution, Depth Upsampling, RAW Burst Alignment, Sparse Neural Rendering, Video Super-Resolution Quality Assessment, Video Saliency Prediction, Blind Photo Quality Assessment.
@@ 6.1 Organizers of AIM 2024 Compressed Video Quality Assessment Challenge
Radu Timofte5,6 (radu.timofte@uni-wuerzburg.de)
@@ 4 MSU Institute for Artificial Intelligence
5 Computer Vision Lab, ETH Zürich, Switzerland
6 Julius Maximilian University of Würzburg, Germany
@@ 6.3 SJTU-MultimediaLab
Ziheng Jia1 (jzhws1@sjtu.edu.cn), Zicheng Zhang1 (zzc1998@sjtu.edu.cn), Wei Sun1 (sunguwei@sjtu.edu.cn), Jiaying Qian1 (2022qjy@sjtu.edu.cn), Yuqin Cao1 (caoyuqin@sjtu.edu.cn), Yinan Sun1 (yinansun@sjtu.edu.cn), Yuxin Zhu1(rye2000@sjtu.edu.cn), Xiongkuo Min1 (minx- iongkuo@sjtu.edu.cn), Guangtao Zhai1 (zhaiguangtao@sjtu.edu.cn)
@@ 6.7 VPT
Wenhui Meng1 (whmeng@whu.edu.cn), Xiaoheng Tan2, Haiqiang Wang2, Xiaozhong Xu2, Shan Liu2, Zhenzhong Chen1
@@
2: Tencent, China
7 Acknowledgements

### v3

@@ 1 Introduction
This challenge is one of the AIM 2024 Workshop  [^1] https://www.cvlai.net/aim/2024/ associated challenges on: sparse neural rendering [40, 41], UHD blind photo quality assessment [20], compressed depth map super-resolution and restoration [11], raw burst alignment [9], efficient video super-resolution for AV1 compressed content [10], video super-resolution quality assessment [38], and video saliency prediction [39].
@@ 6.1 Organizers of AIM 2024 Compressed Video Quality Assessment Challenge
Radu Timofte5 (radu.timofte@uni-wuerzburg.de)
@@ 4 MSU Institute for Artificial Intelligence
5 University of Würzburg, Germany
@@ 6.3 SJTU-MultimediaLab
Ziheng Jia1 (jzhws1@sjtu.edu.cn), Zicheng Zhang1 (zzc1998@sjtu.edu.cn), Wei Sun1 (sunguwei@sjtu.edu.cn), Jiaying Qian1 (2022qjy@sjtu.edu.cn), Yuqin Cao1 (caoyuqin@sjtu.edu.cn), Yinan Sun1 (yinansun@sjtu.edu.cn), Yuxin Zhu1(rye2000@sjtu.edu.cn), Xiongkuo Min1 (minxiongkuo@sjtu.edu.cn), Guangtao Zhai1 (zhaiguangtao@sjtu.edu.cn)
@@ 6.7 VPT
Wenhui Meng1 (whmeng@whu.edu.cn), Zhenzhong Chen1
@@
Acknowledgements
This work was partially supported by the Humboldt Foundation. We thank the AIM 2024 sponsors: Meta Reality Labs, KuaiShou, Huawei, Sony Interactive Entertainment and University of Würzburg (Computer Vision Lab).